### v1

# Topology of the Fermi surface and universality of the metal-metal and metal-insulator transitions: $d$-dimensional Hatsugai-Kohmoto model as an example

Abstract
The earlier theory [1] of the quantum phase transitions related to the change of the Fermi Surface Topology (FST) is advanced. For such transitions the Fermi surface as a quantum critical manifold determined by the Lee-Yang zeros, the order parameter $\mathcal{P}$ as the $d$-volume of the Fermi sea, and the special FST universality class were introduced in [1]. The exactly solvable Hatsugai-Kohmoto (HK) $d$-dimensional ($d=1,2,3$) model of interacting fermions is analyzed. We explore the relation between the Lee-Yang zeros, the Luttinger and the plateau (Oshikawa) theorems. The validity of the Luttinger theorem in the HK model is confirmed. It is shown that the order parameter $\mathcal{P}$ and the FST universality class describe the transitions between metal and band/Mott insulators, as well as the Lifshitz and van Hove gapless-to-gapless transitions. The gapless phases are established to be the Landau Fermi liquids (metals). In addition to the conventional paradigm with a continuous order parameter, we apply the homology theory to analyze the FST transitions. They are critical points of the Morse function. To quantify FST we use the Euler characteristic, which is calculated for each phase of the HK model. We claim that the FST universality class is robust with respect to interactions and other model details, under the condition that the critical points are non-degenerate.
I Introduction
This paper is to the great extend a sequel of our previous work [1], where we studied quantum criticality and quantum transitions between metal and insulating phases, as well as gapless-to-gapless transitions within the metal state of several $d$-dimensional fermionic models ($d=1,2,3$). This was done with the analysis of zeros (or poles) of the partition function. This rigorous theory, pioneered by Yang and Lee [2, *LeeYang:1952], was proven to be very successful for studies of various models at equilibrium and even out of it. For a short list of references, see, e.g., [4, 5, 6, 7, 8, 9, 10, 11], and more references in there. One of the great advantages of this approach is that it can detect a phase transition without any knowledge of the symmetry breaking or the order parameter involved. In case of fermionic or spin models the order accompanying a quantum transition could be quite evasive, like for example, string or brane order parameters, and one needs an extra effort to reveal duality between branes (strings) and conventional (local) orders with broken symmetries [12, 13, 10, 11, 14].
We have shown [1] that the set of the Lee-Yang (LY) zeros of a $d$-dimensional model at $T=0$ defines the (d1)(d-1)-dimensional Fermi surface (FS) which is the manifold of quantum critical points. [^1] We use the term “Lee-Yang zeros” broadly for generic partition function’s zeros expressed via microscopic and thermodynamic parameters (e.g. as in Eq. (9) and (17) below) at finite volume $V$ and/or number of particles $N$, before the thermodynamic limit is taken. In the original work, Yang and Lee [2, *LeeYang:1952] dealt with zeros of the partition function expressed as polynomial of degree $M(V)$ in fugacity. See also [15] for a nice concise presentation. In the proposed terminology, e.g., the Fisher’s zeros are understood as a special case (class) of the LY zeros. The appearance or restructuring of the FS is a quantum phase transition. In the latter case it is called the Lifshitz transition [16]. See also [17, 18, *Volovik:2017, *Volovik:2018] and [21]. Appearance of a (d1)(d-1)-dimensional membrane in the momentum space [21] breaks the initial symmetry of the latter, and as a consequence gapless excitations around the FS appear as well, which is a hallmark of metallic phase. The FS coincides also with the manifold of zeros of the inverse single-particle Green’s function at $T=0$.
To quantify the transition accompanied change of the Fermi Surface Topology (FST) (appearance/disappearance, openning/closure, connectedness change, etc), we introduced in [1] the order parameter $\mathcal{P}$ for the gapless metallic phase via the $d$-volume of the Fermi sea $\mathcal{S}_{d}$. In this paper we consider the $d$-cubic lattice with the first Brillouin zone (BZ) $\mathbf{k}\in[-\pi,\pi]^{d}$, so the definition [1] of the order parameter $\mathcal{P}$ reads:
$\mathcal{P}\coloneqq\mathcal{S}_{d}~\big(\mathrm{mod}(2\pi)^{d}\big)~.$ (1)
Such definition of $\mathcal{P}$ is consistent with the scaling properties of other observables within the special (FST) universality class determined by the critical indices [1]:
$\mathrm{FST:}~~\nu^{\prime}=\nu=\frac{1}{2}~,~~z=2~,~~\beta=d\nu^{\prime}~,~~\gamma=\alpha=1-d\nu^{\prime}~,~~\eta=d~,$ (2)
where all notations are standard [22], except the additional index $\nu^{\prime}$ which is introduced as the critical exponent of the incommensurate (IC) wave number [23]. The FS transitions in two paradigmatic interacting metals – the Fermi and the Tomonaga-Luttinger liquids – belong to the universality class (2). The FST indices satisfy the Fisher identity
$\alpha+2\beta+\gamma=2~,$ (3)
and the (hyper)scaling relations
$\gamma=(2-\eta)\nu~,~~2-\alpha=(d+z)\nu~,~~~2\beta=(d+z-2+\eta)\nu~.$ (4)
The Hamiltonian of interacting spin-1/2 fermions residing on the $d$-cubic lattice we study in this paper, is almost a unique example of the exactly-solvable non-trivial interacting $d$-dimensional model. It has a reach phase diagram with several metallic and insulating phases. It was introduced and solved by Hatsugai and Kohmoto in Ref. [24] (a closely related model was proposed a year before by Baskaran in Ref. [25]). The original study [24] was followed up by several subsequent papers at a moderate pace [26, 27, 28], until recently, when the interest to the Hatsugai-Kohmoto (HK) model has grown considerably, see, e.g., [29, 30, 31, 32, 33, 34] and more references there.
Since the exact solution yields the results for the partition function, LY zeros, Green’s function, etc, via explicit equations at the difficulty level of the statistical mechanics of free fermions, the HK model is an excellent playground to test and to advance the theory [1]. We show in this work that the order parameter (1) and the FST universality class (2) describe the transitions of the HK model from the metal (gapless) phase to the phases of band insulator (BI) or the interaction-driven Mott insulator (MI). The Lifshitz [16] and van Hove [35] transitions between the gapless phases of the model belong to the same class. It appears that the Lifshitz transition in the gapless regime and its critical properties were not addressed in earlier literature on the HK model. It separates two gapless phases (denoted as M and M2) which are established to be conventional Landau Fermi liquids (metals with one and two Fermi surfaces, respectively).
We show in this paper that in addition to the description of the FST transitions within the standard paradigm of critical phenomena (1) and (2), they can be also identified as critical points of the Morse function where the cohomology group changes [36, 37]. The discrete change of the FST is quantified by the Euler characteristic which is calculated for each phase of the HK model.
[…]
The rest of the paper is organized as follows: In Sec. II we introduce the $d$-dimensional Hatsugai-Kohmoto model and give its main exact results. Sec. III contains the results for the phases and transitions of the HK model for $d=1,2,3$. The technical details on the LY formalism, derivations, and the Morse (homology) theory are collected in three Appendices A, B, and C. We encourage readers to go through Appendices A and B before starting Sec. III. The concluding remarks and discussion are given in Sec. IV.
II Model, definitions, and some exact results
Most of equations we collect in this section were either obtained explicitly or follow straightforwardly from the results of the original paper by Hatsugai and Kohmoto [24] and of the following work by Nogueira and Anda [27].
The Hatsugai-Kohmoto (HK) model treats the interacting spin-1/2 fermions hopping on the $d$-dimensional (hyper)-cubic lattice of $L^{d}$ sites. We will analyze the particle-hole symmetric version of the HK Hamiltonian [24]:
$\hat{H}$ $=$ $\sum_{\mathbf{k}}\hat{H}_{\mathbf{k}}~,$
$\hat{H}_{\mathbf{k}}$ $=$ $\varepsilon(\mathbf{k})(\hat{n}_{\mathbf{k}\uparrow}+\hat{n}_{\mathbf{k}\downarrow})+U\Big(\hat{n}_{\mathbf{k}\uparrow}-\frac{1}{2}\Big)\Big(\hat{n}_{\mathbf{k}\downarrow}-\frac{1}{2}\Big)~,$ (5)
$\hat{n}_{\mathbf{k}\sigma}$ $=$ $c^{{\dagger}}_{\mathbf{k}\sigma}c_{\mathbf{k}\sigma},~~\varepsilon(\mathbf{k})=-2t\sum_{i=1}^{d}\cos k_{i}~.$ (6)
The wave vectors are restricted to the first Brillouin zone (BZ) $\mathbf{k}\in[-\pi,\pi]^{d}$. We set the units such that $\hbar=k_{B}=1$, the lattice spacing $a=1$, and the energy scale $2t=1$. Thus
$\varepsilon(\mathbf{k})\in[-d,d]~.$ (7)
In this paper we will analyze the most interesting case of repulsion $U>0$ and $d=1,2,3$.
The eigenvalues of $\hat{H}_{\mathbf{k}}$ are
$\Bigg\{\frac{U}{4};~\varepsilon(\mathbf{k})-\frac{U}{4};~\varepsilon(\mathbf{k})-\frac{U}{4};~2\varepsilon(\mathbf{k})+\frac{U}{4}\Bigg\}~,$ (8)
whence the grand partition function is readily found:
$\mathcal{Z}$ $=$ $\mathrm{Tr}e^{-\beta(\hat{H}-\mu\hat{N})}=\prod_{\mathbf{k}}\mathcal{Z}_{\mathbf{k}}~,$ (9)
$\mathcal{Z}_{\mathbf{k}}$ $=$ $e^{-\beta U/4}+2e^{-\beta(\xi_{\mathbf{k}}-U/4)}+e^{-\beta(2\xi_{\mathbf{k}}+U/4)}~,$ (10)
with $\xi_{\mathbf{k}}\coloneqq\varepsilon(\mathbf{k})-\mu$. The distribution function per spin
$n_{\mathbf{k}\sigma}\coloneqq\langle\hat{n}_{\mathbf{k}\sigma}\rangle=\frac{1}{2}n_{\mathbf{k}}=\frac{1}{2\beta}\frac{\partial\ln\mathcal{Z}_{\mathbf{k}}}{\partial\mu}$ (11)
(note that $n_{\mathbf{k}\uparrow}=n_{\mathbf{k}\downarrow}$) is:
$n_{\mathbf{k}\sigma}$ $=$ $\frac{1+e^{\beta(\xi_{\mathbf{k}}+U/2)}}{1+2e^{\beta(\xi_{\mathbf{k}}+U/2)}+e^{2\beta\xi_{\mathbf{k}}}}$ (12)
$\xrightarrow[\beta\to\infty$]{~} $\frac{1}{2}\big[\Theta(U/2-\xi)+\Theta(-U/2-\xi)\big]$ (13)
It is plotted in Fig. 1 at $T=0$.
Figure 1: (a) The distribution function $n_{\mathbf{k}\sigma}(\xi)$ at $T=0$ given by Eq. (13); (b) The ground-state values $n_{\mathbf{k}\sigma}(\xi)=\{0,\frac{1}{2},1\}$, where $\xi=\varepsilon-\mu$, shown in the plane $(\varepsilon,\mu)$. The bold solid lines are $\mu=\varepsilon\pm U/2$, the energy band $\varepsilon\in[-d,d]$.
The Matsubara single-particle Green’s function for the “particle-hole asymmetric” Hamiltonian $\hat{H}_{\mathbf{k}}^{\prime}=\varepsilon(\mathbf{k})(\hat{n}_{\mathbf{k}\uparrow}+\hat{n}_{\mathbf{k}\downarrow})+U\hat{n}_{\mathbf{k}\uparrow}\hat{n}_{\mathbf{k}\downarrow}$ was found in [27]. For the symmetrized Hamiltonian (5) the Green’s function is [43]
$G_{\sigma}(\mathbf{k},i\omega_{n})=\frac{1-n_{\mathbf{k}\sigma}}{i\omega_{n}-\xi_{\mathbf{k}}+U/2}+\frac{n_{\mathbf{k}\sigma}}{i\omega_{n}-\xi_{\mathbf{k}}-U/2}=\frac{i\omega_{n}-\xi_{\mathbf{k}}+U(n_{\mathbf{k}\sigma}-1/2)}{(i\omega_{n}-\xi_{\mathbf{k}})^{2}-U^{2}/4}~,$ (14)
where $i\omega_{n}=i(2n+1)\pi/\beta$.
To better understand the physical properties of the HK model and its ground state phases, we will be analysing zeros of the partition function (9) in the subsequent sections. From the representation of the $\mathbf{k}$-component of the grand partition function $\mathcal{Z}_{\mathbf{k}}$ (10) as:
$\mathcal{Z}_{\mathbf{k}}=e^{-\beta U/4}\big(e^{-\beta\xi_{\mathbf{k}}}-w_{+}\big)\big(e^{-\beta\xi_{\mathbf{k}}}-w_{-}\big)~,$ (15)
where
$w_{\pm}\coloneqq e^{i(2n+1)\pi}e^{\beta U/2}\big(1\pm\sqrt{1-e^{-\beta U}}\big)~,$ (16)
zeros of the partition function are found:
$\mathcal{Z}_{\mathbf{k}}(\xi)=0~\Longrightarrow~\xi_{\mathbf{k}}=\pm\Bigg(\frac{U}{2}+\frac{1}{\beta}\ln\Big(1+\sqrt{1-e^{-\beta U}}\Big)\Bigg)+i(2n+1)\pi/\beta~.$ (17)
As one can see from the above equation, the zeros lie in the real range ($\xi_{\mathbf{k}}\in\mathds{R}$) only at $T=0$, that is only quantum phase transitions can occur in the HK model in any spatial dimension. [^3] Transitions of the type of disorder line can happen at finite $T$ when the root is complex $\xi_{\mathbf{k}}\in\mathbb{C}$ [10], but we will not discuss this issue in the present work.
We can also find easily poles of the Green’s function (14):
$\xi_{\mathbf{k}}=\pm\frac{U}{2}+i(2n+1)\pi/\beta~.$ (18)
and its zeros:
$\xi_{\mathbf{k}}=U\big(n_{\mathbf{k}\sigma}-1/2\big)+i(2n+1)\pi/\beta~.$ (19)
The Matsubara Green’s function $G_{\sigma}(\xi,0)$ at $T=0$ is plotted in Fig. 2.
Contrary to the non-interacting case [10], zeros of the partition function (17) and poles of the Green’s function (18) are distinct quantities when $U\neq 0$, and they coincide only in the limit $T\to 0$ yielding the single equation:
$\varepsilon(\mathbf{k})-\mu=\pm\frac{U}{2}~.$ (20)
Figure 2: The Matsubara Green’s function $G_{\sigma}(\xi,0)$ (14) at $T=0$. Note that $n_{\mathbf{k}\sigma}(\xi)=1/2$ at $\xi=0$, see Fig. 1a.
Following [1] we use the complex parametrization of the $d$-dimensional wave numbers within the first BZ as:
$z_{i}=e^{\imath k_{i}},~i=1,...,d~,$ (21)
and express the LY zeros at $T=0$ as the roots of the following equation:
$\sum_{i=1}^{d}(z_{i}+z_{i}^{-1})/2=-\mu_{\pm}~,$ (22)
where
$\mu_{\pm}\coloneqq\mu\pm\frac{U}{2}~.$ (23)
We introduce the following formulation of the LT relating the filling $\bar{n}$, i.e., the $d$-volume of the Fermi sea $\mathcal{S}_{d}$, to the LY zeros:
$T=0:~\mathcal{S}_{d}=(2\pi)^{d}\bar{n}=\int_{A_{\scriptscriptstyle LYZ}^{d}}d\mathbf{k}=\Re\int_{C_{\scriptscriptstyle LYZ}^{d}}\prod_{i=1}^{d}\frac{dz_{i}}{\imath z_{i}}$ (24)
In the above definition the first integration is done in the interior of the $d$-dimensional region of the real $\mathbf{k}$-space within the BZ, bounded by the (d1)(d-1)-surface of the LY zeros (20) (Fermi surface). In the second expression for the LT the integration is done in the region of $\mathds{C}^{d}$ space bounded by the solutions of (22). This somewhat terse definition should become clearer when we analyse the model for the cases $d=1,2,3$ in the following sections. We will explicitly show that the LT for the FS holds in the metallic (gapless) phase of the model, and the definition (24) correctly accounts for the filled bands as well. Note that we have no need to use the Green’s function poles or zeros to define the Fermi or Luttinger surfaces for, respectively, the metallic or gapped phases [44, 45]. (See also [46, 47, 48] on related work.) The definition (24) which involves only more primary notion of the LY zeros is universally working for all phases.
III Analysis of transitions: $d=1,2,3$
III.1 What can happen: Band, Lifshitz, Mott, and van Hove transitions
Although the HK model is interacting, its ground-state properties can be understood from analysis of the filling of two (upper/lower) free-like electron bands
$\varepsilon_{\pm}\coloneqq\varepsilon(\mathbf{k})\pm U/2~.$ (25)
Without interaction the model possesses a two-fold degenerate band $\varepsilon(\mathbf{k})$. When the noninteracting chemical potential lies in the interval $\mu_{\circ}\in[-d,d]$, the free model is in the gapless metallic phase. The critical values $\mu_{\circ}=\mp d$ correspond to the transitions into the phase of incompressible BI with empty/fully-filled bands, respectively. The interaction renormalizes these quantum critical points:
$\mu_{\circ}=\mp d~\longmapsto~\mu_{c,1/2}=\mp(d+U/2)\coloneqq\mp\mu_{c}~,$ (26)
while the physics is the same as for free fermions. These band metal-insulator transitions at $\mu=\mu_{c,1/2}$ with critical indices (2) are discussed in [1], and the HK model fits that general framework.
The key role of interaction in the HK model is that it generates two additional quantum phase transitions absent in the free model:
1. In the weakly interacting regime $U<U_{c}$, where
$U_{c}=2d~,$ (27)
the systems stays in the gapless metal state (similarly to the noninteracting case), while the bands are being continuously filled all the way from $\mu=\mu_{c,1}$ up to the point $\mu=\mu_{c,2}$ with both bands totally filled. However, splitting of the bands due to interaction (25) results in two Lifshitz metal-to-metal transitions within the gapless phase at
$\mu_{L,1/2}=\mp(d-U/2)=\mp\frac{1}{2}(U_{c}-U)\coloneqq\mp\mu_{L}~,$ (28)
when appearance/dissapearance of the second Fermi surface occurs.
2. In case of strong interaction $U>U_{c}$, the Lifshitz transition becomes the Mott metal-insulator transition. The splitting of the bands $\varepsilon_{\pm}$ grows with increasing $U$, and the Mott transition occurs at half-filling when the gap opens at
$\mu_{L,1/2}~\mapsto~\mu_{M,2/1}=\pm(U/2-d)=\pm\frac{1}{2}(U-U_{c})\coloneqq\pm\mu_{M}~.$ (29)
The said above can be easily grasped qualitatively from Fig. 3(a,b) where two bands $\varepsilon_{\pm}$ and critical values of the chemical potential in two cases $U<U_{c}$ and $U>U_{c}$ are plotted for d=1d=1.444We used labels (1,2) in the above definitions of the critical points for convenience of the analysis presented along increasing $\mu$ (filling). When there is no risk of confusion we use the simplified notation, e.g., $\mu_{c}$, etc. All $\mu_{\sharp}\geq 0$ by definition.
Although vanishing of the FS at the transition into the band or Mott insulator can be interpreted as its restructuring, and thus, understood as a special case of the Lifshitz transition [16, 18, *Volovik:2017, *Volovik:2018, 1], in this paper we will use the latter term only for the gapless-to-gapless transitions occurring inside the metallic phase, when the FS changes its connectedness.
In general, the quantum transitions correspond to the critical points of the spectrum
$\nabla_{\mathbf{k}}\varepsilon_{\pm}(\mathbf{k})=\nabla_{\mathbf{k}}\varepsilon(\mathbf{k})=0~.$ (30)
All transitions discussed above correspond to the maxima or minima of $\varepsilon(\mathbf{k})$. In $d\geq 2$ the saddle points occur among the critical points of $\varepsilon_{\pm}(\mathbf{k})$. At the saddle (van Hove) points $\mu_{vH}$ the topology of the FS changes, and a phase transition occurs, as to be discussed in the following sections.
Since there is a certain ongoing confusion in the literature regarding fulfillment of the LT in the HK model [24, 30, 32], we will specifically address this point in what follows. The general statement of the LT [38] is that the volume of the Fermi sea does not change by interactions. The LT for the HK model reads in the notations (24):
$2\mathcal{S}_{d}(\mu_{\circ})=\mathcal{S}_{d}(\mu_{+})+\mathcal{S}_{d}(\mu_{-})~.$ (31)
In the pertinent range $|\mu_{\circ}|<d$, Eq. (31) defines the solution $\mu(\mu_{\circ},U)$ as an odd function of $\mu_{\circ}$ for any $U\gtrless U_{c}$. The differential form of the LT (31) is found as:
$\frac{d\mu}{d\mu_{\circ}}=\frac{2\mathcal{D}_{d}(\mu_{\circ})}{\mathcal{D}_{d}(\mu_{+})+\mathcal{D}_{d}(\mu_{-})}~,$ (32)
where we introduced the single-band density of states (DOS) in the standard way:
$\mathcal{D}_{d}(E)\coloneqq\int\delta(E-\varepsilon(\mathbf{k}))\frac{d\mathbf{k}}{(2\pi)^{d}}~.$ (33)
Using the result for compressibility
$\chi=\frac{\partial\bar{n}}{\partial\mu}=\frac{1}{2}\big[\mathcal{D}_{d}(\mu_{+})+\mathcal{D}_{d}(\mu_{-})\big]~,$ (34)
Eq. (32) can be also written as:
$\frac{d\mu}{d\mu_{\circ}}=\frac{\chi_{\circ}(\mu_{\circ})}{\chi(\mu)}~,$ (35)
with the bare susceptibility $\chi_{\circ}(\mu_{\circ})=\frac{\partial\bar{n}}{\partial\mu_{\circ}}$.
Figure 3: (a,b) One-dimensional spectra $\varepsilon_{\pm}(k)$ in the cases $U<U_{c}$ and $U>U_{c}$, respectively. The critical values of chemical potential: (a): ($\mu_{c,1},\mu_{L,1},\mu_{L,2},\mu_{c,2}$), from bottom to the top; (b): ($\mu_{c,1},\mu_{M,1},\mu_{M,2},\mu_{c,2}$), from bottom to the top. (c,d) The interacting chemical potential $\mu$ as a function of $\mu_{\circ}$ in the cases $U<U_{c}$ and $U>U_{c}$, respectively. The cusp of $\mu$ at the Lifshitz point seen in panel (c) becomes a discontinuity at the Mott transition, panel (d). The plots (a) and (c) are done for $U=0.8$, $\mu_{L,\sharp}=\pm 0.6$, $\mu_{c,\sharp}=\pm 1.4$ ; (b) and (d) for $U=3.6$, $\mu_{M,\sharp}=\pm 0.8$, $\mu_{c,\sharp}=\pm 2.8$.
III.2 d=1
III.2.1 Critical properties
The FST critical properties (2) are most pronounced in one spatial dimension, when the phase transitions are of the second kind. Qualitatively, the analysis below bears on some resemblance between the $1d$ HK model and the fermionic ladder dealt with earlier in [1].
The necessary details on the LY formalism for a single band in $1d$ are given in Sec. A.1. We need to take into account two LY zeros for each band of the HK model at $T=0$, cf. (70):
$z_{\sharp\pm}=e^{\pm\imath k_{\sharp}}=-\mu_{\sharp}\pm\imath\sqrt{\mu_{\sharp}^{2}-1}~,$ (36)
where $\mu_{\sharp}$ stands for $\mu_{\pm}$. The LT, cf. Eqs. (24) and (71) yields the occupation number (filling) per spin for all phases ($\forall~U\geq 0$) through a single formula:
$\bar{n}(\mu)=\Re\sum_{\sharp}\int_{z_{\sharp-}(\mu)}^{z_{\sharp+}(\mu)}\frac{dz}{2\pi\imath z}=\frac{1}{2\pi}\Re\{k_{+}+k_{-}\}~,$ (37)
where
$k_{\pm}=\arccos(-\mu_{\mp})~.$ (38)
The spectrum of the HK model allows two types of solutions:
[…]
where $\mathcal{Q}_{\pm}$ admits only two values:
$\mathcal{Q}_{\pm}=\{0,\pi\}~.$ (40)
The case of the real IC roots $k_{\pm}=k_{F\pm}$ corresponds to the quantum critical (gapless, metallic) phase. The Fermi momenta near the critical points behave as:
$k_{F\pm}-\mathcal{Q}_{\sharp}\propto\sqrt{|\mu-\mu_{*}|}~,$ (41)
where $\mu_{*}$ stands for the corresponding critical value of the chemical potential, thus the critical index of the IC wave number $\nu^{\prime}=1/2$.
The gapped phases (BI/MI) correspond to the range where both roots $k_{\pm}$ are complex. The real part of the LY root $\Re k_{\pm}$ gives the commensurate (0,1/2) contribution to the filling, while $\kappa=\min\Im k_{\pm}$ yields the inverse correlation length. Since near the critical points in the gapped phases $\kappa\propto\sqrt{|\mu-\mu_{*}|}$, the critical index $\nu=1/2$, and the equality $\nu^{\prime}=\nu$ holds. The (additive) order parameter (1) of the metallic phase behaves as $\mathcal{P}\propto\sqrt{|\mu-\mu_{*}|}$ near the critical points, yielding $\beta=1/2$. The compressibility
$\chi=\frac{1}{2\pi}\Re\Bigg\{\frac{1}{\sqrt{1-\mu_{+}^{2}}}+\frac{1}{\sqrt{1-\mu_{-}^{2}}}\Bigg\}$ (42)
behaves near the critical points as $\chi\propto|\mu-\mu_{*}|^{-1/2}$, yielding $\gamma=1/2$. The parameters $\bar{n}$ and $\chi$ are shown in Fig. 4. The Mott insulator is incompressible, see Figs. 4(d), and the critical properties of the metallic phase near the Mott transition are determined by the FST universality class (2) with $d=1$.
Upon approaching the critical interaction from both sides $U\to U_{c}\pm 0$, two Lifshitz points $\mu_{L,\sharp}$ or the Mott points $\mu_{M,\sharp}$ merge into a single critical point at $\mu_{L/M,1}\to\mu_{L/M,2}\to 0$, cf. Eqs. (28) and (29). The symmetric (with respect to $\mu\leftrightarrow-\mu$) features seen in Fig. 4 at the Lifshitz and Mott critical points merge at $\mu=0$ when $U\to U_{c}$.
Figure 4: $d=1$ ($U_{c}=2$): (a,b) – filling $\bar{n}(\mu)$; (c,d) – compressibility $\chi(\mu)$ in the cases $U<U_{c}$ and $U>U_{c}$. The plots (a) and (c) are done for $U=0.8$ with the critical values of chemical potential: $\mu_{L,\sharp}=\pm 0.6$, $\mu_{c,\sharp}=\pm 1.4$ . For the plots (b) and (d): $U=3.6$, $\mu_{M,\sharp}=\pm 0.8$, $\mu_{c,\sharp}=\pm 2.8$. The Euler characteristics $\chi_{E}$ for each phase are shown in parentheses, see panels (c,d).
III.2.2 Luttinger and plateau theorems
$\underline{In~the~gapless~regime~(U<U_{c}):}$ the LT (31) amounts to the following equation:
$2\arccos(-\mu_{\circ})=\bar{n}~,$ (43)
where $\mu_{\circ}$ is the chemical potential of the non-interacting model and $\bar{n}$ is given explicitly by the r.h.s. of Eq. (37). This equation we need to solve in order to find $\mu(\mu_{\circ},U)$. The Fermi energy is understood as
$\varepsilon_{F}=\mu(T=0)=\varepsilon(k_{F\pm})\pm U/2~.$ (44)
When the noninteracting chemical potential lies in the interval $\sqrt{1-U/2}<|\mu_{\circ}|<1$, the renormalized potential is bound within the range $\mu_{L}<|\mu|<\mu_{c}$. In this case a simply connected Fermi sea is localized between two Fermi points $\pm k_{F\sharp}$. See Fig. 3 (a) for visualization. When $|\mu_{\circ}|<\sqrt{1-U/2}$, ($|\mu|<\mu_{L}$) the Fermi sea is disconnected with four Fermi points $\pm k_{F\pm}$. Explicitly we find:
$\mathrm{At~~}\sqrt{1-U/2}<|\mu_{\circ}|<1~\Leftrightarrow~\mu_{L}<|\mu|<\mu_{c}:$
$k_{F\sharp}=2\arccos(-\mu_{\circ}),~~\mu=\pm(1-U/2-2\mu_{\circ}^{2})~,$ (45)
and
$\mathrm{At~~}|\mu_{\circ}|<\sqrt{1-U/2}~\Leftrightarrow~|\mu|<\mu_{L}:$
$k_{F\pm}=\arccos(-\mu\pm U/2),~~\mu=\mu_{\circ}\sqrt{1-\frac{U^{2}}{4(1-\mu_{\circ}^{2})}}~.$ (47)
The Lifshitz transition at $\mu=\mu_{L,\sharp}=\pm(1-U/2)$, where the simply connected Fermi sea splits into two disconnected patches, corresponds to $\mu_{\circ}=\pm\sqrt{1-U/2}$. A representative curve $\mu(\mu_{\circ})$ is shown in Fig. 3 (c).
$\underline{In~the~case~U>U_{c}}$, the bands $\varepsilon_{\pm}$ are separated by a gap, see Fig. 3(b). When $\mu$ varies within the range $\mu_{M}<|\mu|<\mu_{c}$, the HK model is in the metal phase with a simply connected Fermi sea localized between $\pm k_{F}=\pm 2\arccos(-\mu_{\circ})$. From invariance of the volume of the Fermi sea we obtain:
$\mathrm{At}~~0<|\mu_{\circ}|<1~\mathrm{(gapless~phase)}:~\mu=(U/2-1+2\mu_{\circ}^{2})~\mathrm{sign}(\mu_{\circ})~,$ (48)
see Fig. 3(d). The gap opens at half-filling when $\mu_{\circ}=0$. The interacting chemical potential (48) is unambiguous when approaching the Mott phase from the metal sides $\mu\to\mp\mu_{M}\mp 0$. However $\mu(\mu_{\circ})$ is an ill-defined function in the Mott phase corresponding to the range $\mu_{M},\mu_{M}]$ with the plateau $\bar{n}=1/2$, seen in Fig. 4 (b). Opening of the gap in the Mott phase appears as a discontinuity of the function $\mu(\mu_{\circ})$, compare Figs. 3 (c) and (d). We redefine [^5] With redefinition (49) the FS (the quantum critical manifold $\Leftrightarrow$ LY zeros $\Leftrightarrow$ poles of the Green’s function) of the noninteracting model at half-filling coincides with the manifold of zeros of the interacting Green’s function, cf. Eq. (19) and Fig. 2. The latter is called Luttinger surface in the literature [44]. Note although the Luttinger surface inherits the property of the FS to preserve its measure, the former, contrary to the latter, does not correspond to the special (singular) points of the interacting partition function (grand potential). The LY formalism proposed in this study is self-sufficient, no need for special points of the Green’s functions to obtain additional results. the function $\mu(\mu_{\circ})$ such that
$\mu(0)\coloneqq 0~.$ (49)
As discussed in Appendix A.1, the satisfaction of the Oshikawa theoremI is guaranteed by the result (71) following from the LY version of the LT. Another way to put it: it is guaranteed by the only two types of solutions (39) for the LY zeros admitted by the model’s spectrum. The gapless metallic phase corresponds to one or two real solutions $k_{\pm}=k_{F\pm}$, cf. Eq. (39) ($\kappa=0$, infinite correlation length). The complex LY roots $k_{\pm}\in\mathds{C}$ corresponding to the gapped phases ($\kappa\neq 0$ $\Leftrightarrow$ finite correlation length) occur with $\Re k_{\pm}=\mathcal{Q}_{\pm}\propto\bar{n}$. Since $\mathcal{Q}_{\pm}$ only admits the values 0 or $\pi$, the filling in the BI/MI phase demonstrates plateaux $\bar{n}=0,1/2,1$ when $(\mathcal{Q}_{-},\mathcal{Q}_{+})=(0,0)$, $(\pi,0)$, $(\pi,\pi)$, respectively. See Figs. 4 (a,b).
Along with the order parameter in the metal phase $\mathcal{P}$ given by the measure of the Fermi sea, on can readily calculate the Euler characteristic which signals the topological transformations of the FS (change of homotopy classes, cohomology groups). See Appendix B for definitions and details. The $1d$ case with two bands can be easily incorporated into the Morse theory outlined in App. B, if we deal with the problem of filling two circles placed at the different ground levels $\varepsilon=-1\mp U/2$. For the gapless case we find using Eqs. (98),(103), and (104):
[…]
where M(M2) denotes metallic phase with a simply connected (disconnected) FS, respectively. $\chi_{E}$ changes at the critical points of band transitions where FS appears/dissapears, as well at the Lifshitz points where the FS changes its connectedness. For the gapped case the Euler numbers are:
[…]
For reader’s convenience we have shown $\chi_{E}$ in different phases directly in Figs. 4 (c,d).
III.3 d=2
III.3.1 Critical properties
The key points of the analysis of this case will be essentially the same as for $d=1$, so we shall be more brief from now. The $2d$ FST transitions are weaker, cf. Eqs. (2). All the technicalities on the LY formalism and the topological properties of critical points are collected in the Appendices A.2 and B. The main difference of $2d$ is the additional transitions at special fillings (van Hove singularities) coming from the saddle points at a half-filled single band. It is insightful to analyze the BI, MI, Lifshitz, and the van Hove transitions of the $2d$ model in the framework of the Morse theory (see Appendix B). These transitions are the critical points of the Morse (height) function which controls the filling of two “collapsed” tori $\mathds{T}^{2}$ with removed central points (see Fig. 8). The communicating tori are placed at the different ground levels $\varepsilon=-2\mp U/2$. The transitions correspond to the critical points when $n$-cells are consecutively glued to the filled manifold (as explained for a single torus in Appendix B) and the manifold’s topology changes, until both tori are fully filled.
We find the filling analytically as:
$\bar{n}=\frac{1}{2}\sum_{\pm}\int_{\max\{-\mu_{\pm},-2\}}^{2}\mathcal{D}_{2}(\varepsilon)d\varepsilon~,$ (52)
where the integration of the DOS (76) yields the Meijer G-function (77). The $2d$ compressibility is found from (34). The representative plots of $\bar{n}$ and $\chi$ are shown in Fig. 5. Qualitatively, the behavior of these parameters at the critical points of the band ($\mu=\pm(2+U/2)$) or Mott/Lifshitz ($\mu=\pm(2-U/2)$) transitions is similar to the $1d$ case discussed in great detail above. The critical indices (2) signal weaker singularities, e.g., the compressibility demonstrates only the cusps at those points. The new $2d$-feature is the van Hove transition at
$\pm\mu_{vH}=\pm U/2~.$ (53)
with logarithmic divergencies in $\chi\propto D_{2}(\epsilon)\propto-\ln|\mu-\mu_{vH}|$, which however belong to the same FST universality class (2) with $\gamma=0$ as the other $2d$ critical points.
Figure 5: $d=2$ ($U_{c}=4$): (a,b) – filling $\bar{n}(\mu)$; (c,d) – compressibility $\chi(\mu)$ in the cases $U<U_{c}$ and $U>U_{c}$. The plots (a) and (c) are done for $U=2.6$ with the critical values of chemical potential: $\mu_{L,\sharp}=\pm 0.7$, $\mu_{c,\sharp}=\pm 3.3$. The van Hove critical points correspond to $\mu_{vH,\sharp}=\pm 1.3$ and $\bar{n}=1/4,3/4$. For the plots (b) and (d): $U=8$, $\mu_{M,\sharp}=\pm 2$, $\mu_{c,\sharp}=\pm 6$, $\mu_{vH,\sharp}=\pm 4$. The Euler characteristics $\chi_{E}$ for each phase are shown in parentheses, see panels (c,d).
The rest of the results we present for two regimes separately:
1). In the gapless regime ($U<U_{c}$) we infer from Eqs. (52), (34), and equations given in Appendix A.2 that in the metal phase $\mu_{c,1}<\mu<\mu_{c,2}$ the model undergoes two Lifshitz and two van Hove transitions, corresponding to the points where the FS changes its topology: simply connected $\leftrightarrow$ disconnected (Lifshitz) or closed (particle-like) $\leftrightarrow$ open (hole-like) (van Hove). Those transitions are clearly seen in Fig. 5 c, plotted for $U=2.6$. The additive order parameter measuring contributions to the Fermi sea from both bands
$\mathcal{P}=\mathcal{P}_{+}+\mathcal{P}_{-}~,$ (54)
[…]
Another clean-cut indication of the transitions where the FS undergoes topological transformations is provided by the Euler characteristic $\chi_{E}$. It changes at the critical points. $\chi_{E}$ is calculated as $\chi_{E}=\chi_{E}^{+}+\chi_{E}^{-}$ (cf. Eq. (104)), where the contributions $\chi_{E}^{\pm}$ of the upper/lower bands are found from (95). For each phase $\chi_{E}$ is shown in Fig. 5 (c).
Note that the order of the Lifshitz and van Hove transitions changes at weak interaction $U<2$: upon increasing filling ($\mu$) the overlap of the lower and upper bands at $\mu=\mu_{L,1}=-2+U/2$ occurs before the lower band becomes half-filled at $\mu=\mu_{vH,1}=-U/2$, cf. (28), similarly on the positive side $\mu>0$. A simple exercise allows to check that the physics does not change, so below we present the results for the case when the van Hove transition occurs in the M1-phase with a single FS ($U>2$). The opposite case is analysed in Appendix C.
2) In the Mott regime ($U>U_{c}$) the bands $\varepsilon_{\pm}$ are separated by the gap, their filling occurs consecutively, there is only one FS for each metal phase, and no Lifshitz transitions exist in this case. Instead, the MI phase appears at $\bar{n}=1/2$ when $\mu_{M}<\mu<\mu_{M}$. The analytical properties of the order parameter $\mathcal{P}$ near the critical points are the same as in the gapless case above. Near at the boundaries with the band/Mott insulator $\mathcal{P}_{\pm}\propto\epsilon$, and $\mathcal{P}_{\pm}\propto-\epsilon\ln\epsilon^{2}$ near the van Hove critical points. Representative plots are given in Figs. 5 (b) and (d). Qualitatively, the gapless phase M2 with two FS is replaced by the incompressible Mott phase. The Euler characteristics for this case are shown in Fig. 5 (d).
III.3.2 Luttinger theorem
The LT (31) in the pertinent range $|\mu_{\circ}|<2$ (cf. Eqs. (83) and (84)) amounts to:
$2g(\mu_{\circ})=g_{R}(\mu_{+})+g_{R}(\mu_{-})~,$ (55)
with the solution $\mu(\mu_{\circ},U)$ as an odd function of $\mu_{\circ}$ for any $U\gtrless U_{c}$.
In the gapless regime, Eq. (55) yields a unique continuous solution $\mu(\mu_{\circ},U)$ for each value of $\mu_{\circ}$ in the whole range $2<\mu_{\circ}<2$. The function $\mu(\mu_{\circ},U)$ found numerically, is qualitatively similar to the one shown in Fig. 3(c) for the $1d$ case, even smoother, because of the weaker $2d$ singularities. Although the function $\mu(\mu_{\circ},U)$ is available only numerically, it is possible to find $\mu$ analytically near special points to leading order, using the results collected in Appendix A.2. The potential $\mu(\mu_{\circ})$ vanishes at zero $\mu_{\circ}$ as
$\mu(\mu_{\circ},U)\approx-\frac{2}{g^{\prime}(U/2)}\mu_{\circ}\ln\frac{\mu_{\circ}^{2}}{4},~~|\mu_{\circ}|\ll 1~.$ (56)
In the vicinity of the BI we find from (79):
$\mu_{\circ}\to\pm 2\mp 0:~\mu=\mp(2-U/2)+2\mu_{\circ}+\mathcal{O}(U^{2})~,$ (57)
valid for $U\gtrless U_{c}$.
In the Mott regime the situation is similar to the $1d$ case shown in Fig. 3(d): a unique continuous solution $\mu(\mu_{\circ},U)$ of Eq. (55) exists in the range $0<|\mu_{\circ}|<2$, but the interacting chemical potential is discontinuous at half-filling:
$\mu_{\circ}\to\pm 0:~~\mu\to\pm\mu_{M}~.$ (58)
To lowest order we find
$\mu\approx\pm(U/2-2)-\frac{2}{\pi}\mu_{\circ}\ln\frac{\mu_{\circ}^{2}}{4}~,~~|\mu_{\circ}|\ll 1$ (59)
As follows from the above equation, when $\mu_{\circ}\to\pm 0$, $\mu$ tends to its critical values $\pm\mu_{M}$ with an infinite slope, in agreement with the exact result (112).
The two van Hove critical points at $\mu=\mp U/2$ correspond to the values (81) of the bare chemical potential $\mu\to\mp\mu_{\circ,s}=\mp 0.7198$, when the single non-interacting band is $1/4$ and $3/4$-filled. Then, correspondingly, the lower or upper band of the interacting model is $1/2$-filled. Using in Eq. (55) the expansions (82) and (80), we get in lowest order the transcendental equation for $\mu$ valid for $U>2$:
$g_{1}(\mu_{\circ}\pm\mu_{\circ,s})=-\mu_{\pm}\ln\frac{\mu_{\pm}^{2}}{4}$ (60)
Closely to the critical point when $|\mu_{\pm}|<2/e$ the approximate solution is:
$\mu\to+\mu_{vH}:~\mu\approx U/2-g_{1}(\mu_{\circ}-\mu_{\circ,s})/\ln\Big(\frac{g_{1}(\mu_{\circ}-\mu_{\circ,s})}{4}\Big)^{2}~.$ (61)
The approximation (61) retains the exact properties following from the LT (55) and (112). At the van Hove points $\mu_{\circ}\to\pm\mu_{\circ,s}$ ($\mu\to\pm U/2$), the interacting chemical potential $\mu(\mu_{\circ})$ is a continuous function without cusps, its slope smoothly vanishes $\mu^{\prime}(\pm\mu_{\circ,s})=0$, and the higher-order derivatives $\mu^{(n)}(\mu_{\circ,s}),~n\geq 2$, are divergent, in agreement with the exact result (112). [^6] The exact Eq. (112) is valid for the M-phase with a single FS, where only one of two terms on the r.h.s. of Eq. (55) contributes. In the M2-phase with two FSs, both terms contribute to $\frac{d\mu}{d\mu_{\circ}}$ in the region(s) where $\mu^{2}_{-}<4]\cap[\mu^{2}_{+}<4]\neq\emptyset$, so the denominator on the r.h.s. of (112) becomes $\mathbf{K}(1-\mu_{+}^{2}/4)+\mathbf{K}(1-\mu_{-}^{2}/4)$. In particular, this recovers the noninteracting limit $\frac{d\mu}{d\mu_{\circ}}=1$ at $U\to 0$. See Eqs. (111) and (112) in Appendix C. In Appendix C we show that the above results hold with inessential numerical differences for the case $U<2$, when the van Hove transition occurs inside the M2 phase, as well as in the limit $U\to 2$, when the Lifshitz and van Hove critical points merge.
Near the Lifshitz transition we find:
[…]
where the bare potential at the Lifshitz point is determined from the following equation:
$g(\mu_{\circ,\scriptscriptstyle L})=\pi^{2}+\frac{1}{2}g(2-U)~.$ (63)
III.4 d=3
This case is not essentially very different from $d=2$ of the previous subsection. Avoiding being too repetitive, we highlight below the key qualitative points without going into technical details. The $3d$ spectrum yields four the van Hove critical points, see sections A.3 and B in Appendices, and not just two as for $d=2$. We introduce an extra label to distinguish the type of the saddle point (S1 and S2, cf. Eq. (99)) and define
$\mu_{\scriptscriptstyle vH}^{\scriptscriptstyle S1}=-1\pm U/2~,~~\mu_{\scriptscriptstyle vH}^{\scriptscriptstyle S2}=1\pm U/2~.$ (64)
The band filling given by the following equation
$\bar{n}(\mu)=\frac{1}{2}\int_{-3}^{3}dx\mathcal{D}_{3}(x)\big(\Theta(\mu_{+}-x)+\Theta(\mu_{-}-x)\big)~,$ (65)
cf. Eqs. (86) and (87), behaves similarly to that shown in Fig. 5 a,b. For BI ($|\mu|>\mu_{c}$), $\bar{n}(\mu)=0,1$. In the gapless regime $U<U_{c}=6$, $\bar{n}(\mu)$ is a continuous monotonous function, as in Fig. 5 a. When $U>U_{c}$, the filling has a plateau $\bar{n}(\mu)=1/2$ in the Mott phase ($|\mu|<\mu_{M}$), as in Fig. 5 b. The Lifshitz or van Hove points are not pronounced in $\bar{n}(\mu)$, while they are well noticeable in the behavior of the compressibility or DOS, cf. Eq. (34). Several characteristic cases of $\chi(\mu)$ at different $U$ are presented in Fig. 6. Similarly to the $2d$ case, the order of appearance of the van Hove and Lifshitz points in the gapless regime can change, depending on the interaction strength, cf. Eqs. (28) and (64). At two particular values $U=2,4$ two of the van Hove points and the Lifshitz points merge. According to (90), the order parameter (1) vanishes at the critical points $\mu_{c},\mu_{M}$ of transitions into the BI or MI phases with exponent $\beta=3/2$. It is non-vanishing and continuous, but it has an additive contribution with the same exponent at the metal-to-metal (Lifshitz, van Hove) transitions. The compressibility clearly demonstrates those non-analyticities with $\gamma=-1/2$ at the critical points. The topology of the FS surface changes at each critical point, which is signalled by change of the Euler characteristic $\chi_{E}$. The latter is calculated using Eqs. (64),(99),(100), and (104). It is shown in Fig. 6.
Figure 6: $d=3$ ($U_{c}=6$): Compressibility $\chi(\mu)$ for different interaction strengths. (a) $U=1$, critical values of chemical potential are: $\mu_{c,\sharp}=\pm 3.5$, $\mu_{L,\sharp}=\pm 2.5$, $\mu_{vH,\sharp}=\pm 1.5,\pm 0.5$. (b) $U=5$, $\mu_{c,\sharp}=\pm 5.5$, $\mu_{L,\sharp}=\pm 0.5$, $\mu_{vH,\sharp}=\pm 3.5,\pm 1.5$. (c) Two insets for the special cases when the Lifshitz and van Hove points merge. Upper left: $U=2$, $\mu_{c,\sharp}=\pm 4$, $\mu_{L,\sharp}=\mu_{vH,1/4}=\pm 2$, $\mu_{vH,2/3}=0$. Lower right: $U=4$, $\mu_{c,\sharp}=\pm 5$, $\mu_{L,\sharp}=\mu_{vH,2/3}=\pm 1$, $\mu_{vH,1/4}=\pm 3$. (d) $U=10$, $\mu_{c,\sharp}=\pm 8$, $\mu_{M,\sharp}=\pm 2$, $\mu_{vH,\sharp}=\pm 6,\pm 4$. The Euler characteristics $\chi_{E}$ for each phase are shown in parentheses.
The interacting chemical potential $\mu(\mu_{\circ},U)$ is found numerically from the LT (31) and (86). It agrees with (26), i.e., $\mu(\pm 3,U)=\pm(3+U/2)$. At $U<U_{c}$, $\mu(\mu_{\circ})$ is a monotonous odd function, similar to the one shown in Fig. 3 (c), with $\mu(0,U)=0$. In the gapped regime $U>U_{c}$, $\mu(\mu_{\circ})$ undergoes discontinuity in the Mott phase at half-filling ($\mu_{\circ}=0$), similarly to Fig. 3 (d):
$\mu_{\circ}\to\pm 0:~\mu\to\pm(U/2-3)~.$ (66)
IV Conclusion and discussion
The earlier theory of the FS-transitions [1] is corroborated and advanced further. The exactly solvable $d$-dimensional HK model allows in particular to elucidate the role of interactions. The model demonstrates transitions accompanied by the appearance/disappearance of the FS between metal and the band or interaction-driven Mott insulators. It also undergoes the Lifshitz and van Hove gapless-to-gapless transitions in the metal phases, related to the topological changes of the FS (connectedness, openness, closure). All these transitions are described by the non-local topological order parameter (1) and the FST universality class (2). At the transitions into the band/Mott insulator the order parameter (1) vanishes; and it demonstrates cusps at the metal-metal (Lifshitz, van Hove) transitions.
In addition to the standard paradigm of critical phenomena (1) and (2), we apply the homology (Morse) theory to analyze the metal-metal and metal-insulator transitions. They are identified as critical points of the Morse function where the cohomology group, the Betti numbers, and some other topological quantities of the filled Fermi sea change. To quantify discrete changes of the FS topology we use the Euler characteristic, which is calculated for each phase.
We build up the theory with the LY zeros as the primary fundamental point of the analysis, since the partition function gives directly the grand potential and other thermodynamic parameters. From the exact results of the HK model we explore the relation between the LY zeros, the Luttinger and the plateau (Oshikawa) theorems. In particular, we introduced the general (model-independent) formulation of the Luttinger theorem yielding the filling $\bar{n}$ via the integration in the region of $\mathds{C}^{d}$ complex space bounded by the LY zeros (24). This formulation relating $\bar{n}$ (or the $d$-volume of the Fermi sea $\mathcal{S}_{d}$) to the LY zeros is shown to be working in the metal (gapless) phases, it correctly accounts for the filled bands, and it applies continuously to the gapped phases as well.
The general statement of the Luttinger theorem that the volume of the Fermi sea does not change by interactions, is shown to hold in the HK model via explicit calculations for $d=1,2,3$. The analytic expressions (31), (32) of this theorem for the HK model determine the interacting chemical potential $\mu(\mu_{\circ},U)$ in the gapless regime ($U<U_{c}$) as a continuous odd function of $\mu_{\circ}$. The latter is the chemical potential of the non-interacting model. Several explicit formulas for $\mu(\mu_{\circ},U)$ are calculated near the critical points (band, Lifshitz, Mott, van Hove). In the regime of strong interaction ($U>U_{c}$) the Mott transition occurs with the gap opening at half-filling, when $\mu_{\circ}=0$. A discontinuity is found in the interacting chemical potential $\mu(\mu_{\circ}\to\pm 0)\to\pm\mu_{M}$.
With redefinition $\mu(0)\coloneqq 0$ at half-filling, the FS as a quantum critical manifold determined by the LY zeros, coincides with the manifold of zeros of the interacting Green’s function (19), called the Luttinger surface [44, 45]. Although it inherits the property of the FS to preserve its measure, the former, contrary to the latter, does not correspond to the special (singular) points of the interacting partition function. As seen from the exact results of the HK model, the poles/zeros of Green’s function do not coincide with the LY zeros in general ($U\neq 0$, $T\neq 0$). On the other hand, the LY formalism is self-sufficient, and we do not need to invoke Green’s function to deal with the FS, commensurate or IC fillings, and plateaux.
The gapless phases (M, M2) of the HK model are characterized by:
(i) interaction-independent $d$-volume of the Fermi sea;
(ii) one (M) or two (M2) Fermi surfaces defined by $\mu(\mu_{\circ},U)$;
(iii) simple poles (18) of the Green’s function.
So, these phases are conventional Landau Fermi liquids (metals).
The important question to address: how robust and universal the FST universality class (2) is? Whatever is the model Hamiltonian of a metal, [^7] It is not even necessary to be a normal metal, i.e., the Fermi liquid. In the XXZ chain in the external field $h$, the Fermi sea (order parameter) in the gapless Tomonaga-Luttinger phase vanishes near the critical point $h_{c}$ as $\sqrt{h_{c}-h}$ [52], in agreement with (2). it suffices to have a conduction band available for filling. If we assume the band energy function $\varepsilon(\mathbf{k})$ to be bounded and differentiable at least three times, then the result (2) should hold. The periodic function $\varepsilon(\mathbf{k})$ can be identified with the Morse function. For different models their Morse functions, the number of critical points, their order of appearance and multiplicity are different. But there are only maxima, minima, and saddle points (for $d\geq 2$) among the critical points of any Morse function. The critical points are specified by $(d,s)$: for a given $d$ the index $s$ can admit values in the range $0\leq s\leq d$. The maxima and minima correspond to the band, Mott, or Lifshitz transitions, while the saddle points correspond to the van Hove transitions. Only the signature of the expansion of the Morse function near the critical point (determined solely by $d$ and $s$) matters for the values of the critical exponents $\beta$, $\gamma$, giving the FST universality (2).
Along with the universal behavior of the order parameter, compressibility (density of states), etc, characterized by the exponents (2), at the transition points topology of the FS undergoes discrete transformations. The type and sequence of those transformations are not universal and depend on the model’s details. For instance for the model with $d$-cubic lattice the Morse function is defined on the $d$-dimensional torus $\mathds{T}^{d}$. In this work we considered several examples of how the topology of the FS undergoes consecutive transformation when $n$-cell (or cells) are added at each critical point, which is signalled by a change of the Euler characteristic, until the band gets fully filled and $\mathrm{FS}\mapsto\mathds{T}^{d}$.
At the degenerate critical points where the Hessian of $\varepsilon(\mathbf{k})$ vanishes, the Morse theory is not applicable. Such high-order critical points occur at some particular (fine-tuned) values of the microscopic parameters, and we assume the situation with non-degenerate critical points of the Morse function as generic. The critical exponents at the high-order points are different from (2), see, e.g., [49, 50, 51] and need to be studied on the case-by-case basis.
Acknowledgements.
I thank A.M. Povolotsky for useful discussions and A.-M. Tremblay for bringing important references to my attention. I am very grateful to V.A. Timorin for reading the manuscript and illuminating discussions of the Morse theory. The author acknowledges financial support from the Russian Science Foundation (RSF), grant No. 24-22-00075.
Appendix A Details on the Lee-Yang formalism and other derivations
To give necessary technical details on the results used in the main text, we consider a single noninteracting band. We analyze the solutions of
$\sum_{i=1}^{d}\cos k_{i}=\varepsilon~,$ (67)
in the complex parametrization (21) with $\varepsilon\in\mathds{R}$, $\infty<\varepsilon<\infty$ for the three cases $d=1,2,3$. The critical points of the spectrum (30) correspond to the following equations:
$\frac{\partial\varepsilon}{\partial\ln z_{i}}=-\frac{1}{2}(z_{i}-z_{i}^{-1})=0~,~~i=1,...,d.$ (68)
A.1 d=1
In this case the roots of Eqs. (67) and (21) are:
$z_{\pm}=-\varepsilon\pm\sqrt{\varepsilon^{2}-1}~,~~z_{+}z_{-}=1~.$ (69)
We trace their path on the complex plane $z$ as $\varepsilon$ varies from -$\infty$ to $\infty$. The plane has a branch cut along the negative real semi-axis, see Fig. 7. $z_{\pm}$ are both real positive/negative for $\varepsilon<-1$ ($\varepsilon>1$), respectively, and they are unimodular complex conjugate numbers at $|\varepsilon|\leq 1$.
$z_{+}=z_{-}^{\ast}~,~~z_{+}=e^{\imath k_{F}}=-\varepsilon+\imath\sqrt{1-\varepsilon^{2}}~.$ (70)
Figure 7: Path of the LY zeros $z_{\pm}(\varepsilon)$ on the complex plane $z$ as $\varepsilon$ varies from -$\infty$ to $\infty$. The roots $z_{\pm}$ on the upper/lower edges of the branch cut along the negative real semi-axis are defined such that $\arg z_{\pm}=\pm\pi$.
The complex version of the LT (24) yields:
$\bar{n}(\varepsilon)=\Re\int_{z_{-}(\varepsilon)}^{z_{+}(\varepsilon)}\frac{dz}{2\pi\imath z}=\Re\frac{1}{2\pi\imath}[\ln z_{+}(\varepsilon)-\ln z_{-}(\varepsilon)]=\frac{\arg z_{+}(\varepsilon)}{\pi}$ (71)
As one can see from the above result, the formulation of the LT via the LY zeros works smoothly for both gapped and gapless phases. Eq. (71) connects $\bar{n}$ to the measure of the Fermi sea $2\pi\bar{n}=2k_{F}=2\arccos(-\varepsilon)$ for the partially filled band ($|\varepsilon|<1$), and it works for the empty and fully filled bands as well, $\bar{n}=0,1$ ($|\varepsilon|>1$).
The LT (71) immediately leads to the LSM-type plateau theorem by Oshikawa [41]: in the gapped phases when $|z_{\pm}|\neq 1$, the spectrum yields $\arg z_{\pm}=0,\pm\pi$, so only commensurate fillings $\bar{n}=0,1$ can occur. For the gapless phase the condition of criticality $|z_{\pm}|=1$ must hold, and $\arg z_{+}=k_{F}$ yields the incommensurate filling.
Two extrema (min/max) of the spectrum (68) at $z_{\pm}=-\varepsilon=\mp 1$ correspond to the quantum phase transitions between the metal and BI. At $\varepsilon=-1$ the two roots merge at the real axis with $z_{+}=z_{-}=1$, see Fig. 7. At $\varepsilon=1$ the roots tend to $z_{+}=z_{-}=-1$ from the opposite edges of the branch cut: $z_{\pm}=\exp(\pm\imath\pi)$. The critical points $\varepsilon=\pm 1$ bound two unit semicircles on the complex plane, which are the arcs of quantum criticality corresponding to $1<\varepsilon<1$, see Fig. 7.
A.2 d=2
The equations (68) for the critical points yield the minimum $\varepsilon=-2$ at $z_{x}=z_{y}=1$, the maximum $\varepsilon=2$ at $z_{x}=z_{y}=-1$, and the saddle (van Hove) points $\varepsilon=0$ at $z_{x}=-z_{y}=\pm 1$ when the band is half-filled. In the gapless range $1<\varepsilon<1$, the FS is determined by the LY roots where $|z_{x}|=|z_{y}|=1$. From the LT (24) one can get the filling as a single integral over a real parameter:
$(2\pi)^{2}\bar{n}(\varepsilon)=2\Im\int_{0}^{\pi}dk[\ln z_{+}(\varepsilon,k)-\ln z_{-}(\varepsilon,k)]=\mathcal{S}_{2}(\varepsilon)~,$ (72)
where
$z_{\pm}(\varepsilon,k)=-a\pm\sqrt{a^{2}-1}~,~~a\coloneqq\varepsilon+\cos k~,$ (73)
and $\mathcal{S}_{2}(\varepsilon)$ denotes the area of the reciprocal space filled by the levels $\varepsilon(\mathbf{k})=-\cos k_{x}-\cos k_{y}\leq\varepsilon$. Note that $z_{+}z_{-}=1$. We chose a branch cut along the negative real semi-axis of $z$, and the roots $z_{\pm}$ on the upper/lower edges of the branch cut are defined such that $\arg z_{\pm}=\pm\pi$.
As in the $1d$ case, analysis of the positions of $z_{\pm}$ on the complex plane as $k$ varies from 0 to $\pi$ at different $\varepsilon\in(-\infty,+\infty)$, allows us to get a better insight on the connection between the LY zeros, Luttinger’s and Oshikawa’s theorems. In the gapped states when $|\varepsilon|>2$, both roots $z_{\sharp}(\varepsilon,k)\in\mathds{R}$. For $\varepsilon\lessgtr\mp 2$, the roots $z_{\sharp}(\varepsilon,k)\gtrless 0$, $\forall k\in[0,\pi]$. Then the integration in (72) is trivial, yielding
[…]
In the gapless region $|\varepsilon|<2$, $z_{\pm}$ are unimodular complex conjugate numbers $z_{+}=z_{-}^{\ast}$ in certain ranges of $k$. At $2<\varepsilon<0$, only $z_{\pm}$ sitting on the arcs of the unit circle $|z|=1$ contribute to (72), resulting in an IC filling. At the saddle critical point $\varepsilon=0$, when the diamond-shaped FS touches the edges of the BZ [53], the roots $z_{\pm}$ run across whole upper/lower semicircles during the integration in (72). At $\varepsilon>0$, $\bar{n}(\varepsilon)$ acquires additional contributions from the parts where $z_{\pm}$ are located on the negative real semi-axis.
It is convenient to bring (72) to the following form:
$\mathcal{S}_{2}(\varepsilon)=4\int_{\max\{-\varepsilon,-2\}}^{2}\mathbf{K}(1-x^{2}/4)dx~,$ (75)
where the complete elliptic integral of the first kind $\mathbf{K}$ in the integrand comes from the DOS. The latter is known from the literature, see, e.g., [53]:
$\mathcal{D}_{2}(\varepsilon)=\int\delta(\varepsilon-\varepsilon(\mathbf{k}))\frac{d\mathbf{k}}{(2\pi)^{2}}=\frac{1}{\pi^{2}}\mathbf{K}(1-\varepsilon^{2}/4)\Theta(4-\varepsilon^{2})~,$ (76)
where $\Theta(x)$ is the Heaviside step function. In the gapless phase we find
$\mathcal{S}_{2}(\varepsilon)=4\int_{-\varepsilon}^{2}\mathbf{K}(1-x^{2}/4)dx=2\pi^{2}+\frac{\varepsilon}{\pi}\mathcal{G}(\varepsilon^{2}/4)~,~~|\varepsilon|\leq 2~,$ (77)
where $\mathcal{G}(x)$ is Meijer G-function. [^8] This is the Mathematica built-in function $MeijerG[\{\{1/2,1/2,1/2\},\{\}\},\{\{0,0\},\{-1/2\}\},x^{2}/4]$. It is convenient to introduce an odd auxiliary function
$g(x)\coloneqq\frac{x}{\pi}\mathcal{G}(x^{2}/4)~.$ (78)
Near the band end the Taylor expansion of $\mathcal{G}$ yields:
$g(x)=2\pi^{2}-2\pi(2-x)-\frac{\pi}{4}(2-x)^{2}-\frac{5\pi}{96}(2-x)^{3}+\mathcal{O}((2-x)^{4}),~~x\sim 2^{-}~,$ (79)
and near half-filling:
$g(x)=-2x\ln\frac{x^{2}}{4}+\mathcal{O}(x),~x\sim 0.$ (80)
For the analysis of the interacting case we will need the expansion of $g(x)$ near the special points $\mp\zeta_{\circ}$ where the band is $1/4$- and $3/4$-filled:
$\zeta_{\circ}=0.71984482...:~~g(\pm\zeta_{\circ})=\pm\pi^{2}~,$ (81)
and $g(x)$ is analytical:
$g(x)=\pi^{2}+g_{1}(x-\zeta_{\circ})-g_{2}(x-\zeta_{\circ})^{2}+g_{3}(x-\zeta_{\circ})^{3}+\mathcal{O}((x-\zeta_{\circ})^{4})~.$ (82)
Numerically, $g_{1}=4\mathbf{K}(1-\zeta_{\circ}^{2}/4)=9.82789$, $g_{2}=2.58482$, and $g_{3}=1.31111$.
With the function $g_{R}(x)$ defined as:
[…]
the filled area is
$\mathcal{S}_{2}(\varepsilon)=2\pi^{2}+g_{R}(\varepsilon)~,~~\forall~\varepsilon\in(-\infty,\infty)~.$ (84)
Note also that
$g_{R}^{\prime}(\varepsilon)=4\pi^{2}\mathcal{D}_{2}(\varepsilon)~,$ (85)
as follows from definitions (83) and (76).
A.3 d=3
The equations (68) for the critical points yield the minimum $\varepsilon=-3$ at $(z_{x},z_{y},z_{z})=(1,1,1)$, the maximum $\varepsilon=3$ at $(z_{x},z_{y},z_{z})=(-1,-1,-1)$, and two saddle (van Hove) points with $\varepsilon=-1$ at $(z_{x},z_{y},z_{z})=(1,1,-1)$; (1,1,1)(1,-1,1); (1,1,1)(-1,1,1), and with $\varepsilon=1$ at $(z_{x},z_{y},z_{z})=(1,-1,-1)$; (1,1,1)(-1,1,-1); (1,1,1)(-1,-1,1). Cf. also Eq. (99). In the gapless range $3<\varepsilon<3$, the FS is determined by the LY roots where $|z_{x}|=|z_{y}|=|z_{z}|=1$. The Lee-Yang formulation of the LT (24) can be brought to the form similar to (72) with an extra integration. The key steps of analysis and the outcomes are analogous to those of the case $d=2$. The filled volume is found as follows:
$\mathcal{S}_{3}(\varepsilon)=(2\pi)^{3}\bar{n}(\varepsilon)=(2\pi)^{3}\int_{-3}^{3}dx\mathcal{D}_{3}(x)\Theta(\varepsilon-x)~,$ (86)
where the single-band DOS is
$\mathcal{D}_{3}(\varepsilon)=\frac{1}{\pi^{3}}\int_{\varepsilon-1}^{\varepsilon+1}\frac{dz}{\sqrt{1-(z-\varepsilon)^{2}}}\mathbf{K}(1-z^{2}/4)\Theta(4-z^{2})~.$ (87)
It vanishes near the band ends $\varepsilon=\pm 3$ as:
$\mathcal{D}_{3}(\varepsilon)\propto\sqrt{|\varepsilon\mp 3|},~~\varepsilon\to\pm 3^{\mp}~,$ (88)
and in the vicinity of the van Hove (saddle) points $\varepsilon=\pm 1$ it behaves as:
$\mathcal{D}_{3}(\varepsilon)-\mathcal{D}_{3}(\pm 1)\propto\sqrt{|\varepsilon\mp 3|}\Theta(-1\pm\varepsilon)+\mathcal{O}(\varepsilon\mp 1),~~\varepsilon\to\pm 1~,$ (89)
where $\mathcal{D}_{3}(\pm 1)=0.2893$. Note that the non-analytic square-root term appears only from one side of the van Hove points. All this leads to the non-analytic contributions to the FS volume near the critical points (99):
$\delta\mathcal{S}_{3}(\varepsilon)\propto|\varepsilon-\varepsilon_{c}|^{3/2}~,$ (90)
so the exponents $\gamma$ and $\beta$ agree with the general result (2). Note that for the $3d$ spectrum (67), the van Hove critical point occurs at a generic IC filling $\bar{n}(\mp 1)=0.21327/0.78222$, while the point of half-filling $\varepsilon=0$ is not critical.
Appendix B Quantum transitions and Morse theory
In this Appendix we explore the topological nature of the metal-metal and metal-BI quantum phase transitions, advancing the pioneering ideas of the classical van Hove paper [35]. The key idea is to apply the Morse theory [36] and more generally, the homology theory [37] for the analysis of the band filling. Due to its periodicity, the spectrum $\varepsilon(\mathbf{k})$ can be defined on a $d$-dimensional torus $\mathds{T}^{d}$ and it satisfies the conditions for the Morse function [36]. The mathematical definition of the critical point in the Morse theory $\nabla_{\mathbf{k}}\varepsilon(\mathbf{k})=0$ coincides with the physical condition for the transition (30). The universal nature of the critical points corresponding to the band, Mott, Lifshitz and van Hove transitions can be understood within a single framework of the cohomology theory outlined below.
B.1 Plain vanilla spectrum (67)
We start with the $\mathbf{2d}$ case of (67) which is non-trivial and easy to visualize. The topologically distinct cases of band filling amount to the standard problem of the Morse theory: analysis of the height function for the torus, see Fig. 8 a. It can be simply formulated as a problem of, say, filling with the water a glassy doughnut touching the ground at single point [36].
We define $M^{\varepsilon}$ as a manifold which is the “filled” part of the torus $\mathds{T}^{d}$ at a given $\varepsilon$:
$\forall~\mathbf{k}\in M^{\varepsilon}\subseteq\mathds{T}^{d}:~\varepsilon(\mathbf{k})\leq\varepsilon$ (91)
We will be particularly interested in the $d$-area $\mathcal{S}_{d}(M^{\varepsilon})$ which measures the filled part of the BZ (FS).
Figure 8: (a) Consecutive building up the torus $\mathds{T}^{2}$ by attaching $n$-cells, as discussed in the text. Adapted from [37]. Indices of the critical points (1-4): min (1) $s=0$, saddle points (2,3) $s=1$, and max (4) $s=2$. (b) The collapsed doughnut with a punctured cental point (its inner radius $r_{inn}\to 0$) shown from different angles.
A generic $2d$ torus yields four critical points (1-4 in Fig. 8 a). Points 1 and 4 correspond, respectively, to the minimum and maximum of the Morse function; and points 2 and 3 are the saddle points. At each critical point the topology of the filled part of the torus surface changes. The point 1 corresponds to attaching 0-cell to the empty space; at the saddle point 2, one $1$-cell (handle) is added, and the second handle is added at the saddle point 3. The 2-cell (membrane) is added at the maximum 4. The homotopies of the manifolds between the critical points are different, and so are their cohomology groups [36, 37].
To numerically account for the change of topology at the critical points one can use the Euler characteristic $\chi_{E}$. For the manifold $M$ it can be calculated in two ways [37]:
$\chi_{E}(M)$ $\overset{\forall f}{=}$ $\sum_{s\geq 0}(-1)^{s}C_{s}(f)$ (92)
$=$ $\sum_{n\geq 0}(-1)^{n}b_{n}~,$ (93)
where $C_{s}$ is the number of the critical points with index ss, [^9] The index $s$ is defined as the number of negative quadratic terms in the expansion of the Morse function $f$ in the vicinity of the critical point. In our case $f=\varepsilon(\mathbf{k})$. and $b_{n}$ are the Betty numbers. For $\mathds{T}^{2}:$ $b_{0}=1,b_{1}=2,b_{2}=1$. It is important to emphasize that Eq. (92) holds for any Morse function, and it is determined solely by the parameters of the cohomology group of $M$ as expressed by Eq. (93).
Special symmetries can lead to merging of the saddle points. In our case the $x\leftrightarrow y$ symmetry of the spectrum leads to a single saddle point in the center of the “collapsed” doughnut (i.e. the doughnut with a punctured cental point when its inner radius $r_{inn}\to 0$, see Fig. 8 b). The saddle point is non-degenerate since the Hessian at $(k_{x},k_{y})=(\pi,0)$ and $(0,\pi)$ is equal to 1-1. The critical points are:
$\mathrm{Min}$ $:$ $~~\varepsilon(0,0)=-2,~~s=0,~C_{0}=1,$
$\mathrm{S}$ $:$ $~~\varepsilon(0,\pi)=\varepsilon(\pi,0)=0,~~s=1,~C_{1}=2,$
$\mathrm{Max}$ $:$ $~~\varepsilon(\pi,\pi)=2,~~s=2,~C_{2}=1.$ (94)
Using the above numbers $C_{s}$ (or equivalently, $b_{n}$) which change when $M^{\varepsilon}$ goes through the critical point [36, 37], we find:
[…]
The minimum/maximum ($\varepsilon=\mp 2$) correspond to the appearance/dissapearance of the FS, respectively. The saddle point at $\varepsilon=0$ (called the van Hove singularity in the literature) corresponds to the half-filled band with the diamond-shaped FS which touches the BZ edges by its corners, see, e.g. [53]. After crossing this point the closed particle-like FS becomes an open hole-like one at $\varepsilon>0$. However a shift by the reciprocal lattice vector $\mathbf{Q}=(\pi,\pi)$ with $\varepsilon(\mathbf{k}+\mathbf{Q})=-\varepsilon(\mathbf{k})$ allows us to switch to the hole representation within the new BZ with a closed hole-like FS. The area of the hole Fermi sea (hole order parameter) can be easily found from equation $\mathcal{S}_{2,h}(\varepsilon)|_{\varepsilon>0}=4\pi^{2}-\mathcal{S}_{2}(\varepsilon)|_{\varepsilon<0}$. At the BI critical points $\mathcal{S}_{2}(-2)=\mathcal{S}_{2,h}(2)=0$, and at the van Hove point $\mathcal{S}_{2}(0)=\mathcal{S}_{2,h}(0)=2\pi^{2}$. From the topological point of view, this is a manifestation of the Poincaré duality. Upon changing the sign of the Morse function $f$, the critical point with index $s$ of $f$ becomes the critical point of index dsd-s of f-f, and $C_{s}\mapsto C_{d-s}$ (and $b_{n}\mapsto b_{d-n}$). Thus, $max\leftrightarrow min$, while the saddle points stay.
Since the order parameter (1) is defined modulo the fully filled band, it could be more natural to use the combined particle-hole order parameter, normalized by the area of the half-filled band, which appears to be more informative for the general analysis:
$\mathcal{P}_{2,ph}(\varepsilon)\coloneqq 1-\frac{g_{R}(\varepsilon)}{2\pi^{2}}\mathrm{sign}(\varepsilon)~.$ (96)
It vanishes at the boundaries of the metal phase where the FS disappears, and it signals the topological transformations of the FS in the metallic phase by a cusp. [^10] Introducing the particle-hole parameter in $1d$ is not useful, since the point of half-filling $\varepsilon=0$ is not critical. See Fig. 9.
Figure 9: Two normalized order parameters for the gapless phase, $d=2$: the particle parameter $\mathcal{P}_{2}(\varepsilon)\coloneqq\mathcal{S}_{2}(\varepsilon)/\mathcal{S}_{2}(0)$, where $\mathcal{S}_{2}(\varepsilon)$ is defined by Eq. (84), and the particle-hole parameter $\mathcal{P}_{2,ph}(\varepsilon)$ (96). By definition $\mathcal{P}_{2,ph}(\varepsilon)=\mathcal{P}_{2}(\varepsilon)$ at $\varepsilon<0$.
The $\mathbf{1d}$ case of the circle $\mathds{T}^{1}=\mathds{S}^{1}$ easily fits in the above analysis with a single minimum/maximum at the ends of the interval $\varepsilon=\mp 1$:
$\mathrm{Min}$ $:$ $~~\varepsilon(0)=-1,~~s=0,~C_{0}=1,$
$\mathrm{Max}$ $:$ $~~\varepsilon(\pi)=1,~~s=1,~C_{1}=1.$ (97)
The Euler characteristic is:
[…]
In the case $\mathbf{d=3}$ the Morse function $\varepsilon(k_{x},k_{y},k_{z})$ is defined on the $3d$ torus $\mathds{T}^{3}$. We will not attempt to visualize filling it, like we did for $\mathds{T}^{2}$ in Fig. 8, but proceed analytically getting some help for intuition from Fig. 10. The Morse function on $\mathds{T}^{3}$ has 8 critical points: 1 minimum, 1 maximum, and 2 saddle points (S1, S2) of multiplicity 3 each. $S1:(\pi,0,0)+2$ permutations and $S2:(\pi,\pi,0)+2$ permutations [35]:
$\mathrm{Min}$ $:$ $~~\varepsilon(0,0,0)=-3,~~s=0,~C_{0}=1,$
$\mathrm{S1}$ $:$ $~~\varepsilon(S1)=-1,~~s=1,~C_{1}=3,$
$\mathrm{S2}$ $:$ $~~\varepsilon(S2)=1,~~s=2,~C_{2}=3,$
$\mathrm{Max}$ $:$ $~~\varepsilon(\pi,\pi,\pi)=3,~~s=3,~C_{3}=1.$ (99)
The Betty numbers are $b_{0}=1$, $b_{1}=3$, $b_{2}=3$, $b_{3}=1$. Similarly to the $2d$ case, at the critical points $n$-cells are glued to $M^{\varepsilon}$, so its topology (cohomology group) changes. We quantify this change with the Euler characteristic, which is found as follows:
[…]
As one can see from Fig. 10, the first saddle point ($S1$) at $\varepsilon=-1$ corresponds to the appearance of the FS necks, while at the second point ($S2$) at $\varepsilon=1$ the necks on the faces of the BZ become open.
Figure 10: The Fermi surfaces of the 3D band (67) at different fillings, shown within the cube of the first BZ $\pi,\pi]^{3}$. The first saddle point ($S1$) at $\varepsilon=-1$ (b), when the vertices of the FS touch the faces of the BZ, precedes appearance of a neck, i.e., closed FS on the faces of the BZ, (c,d). After crossing the second saddle point ($S2$) at $\varepsilon=1$ (e), the necks on the faces become open surfaces, (f).
The class of universality of all topological transitions at the critical points discussed in this subsection is determined by the critical exponents (2). As was first done by van Hove for $d=2,3$ [35], the index $\gamma$, for instance, can be extracted from the DOS calculated in the vicinity of critical points. It is determined solely by the topology of $\mathds{T}^{d}$: only the signature of the quadratic expansion of the Morse function defined on $\mathds{T}^{d}$ (i.e., the index $s$) and the spatial dimension $d$ matter to get the result (2).
The $1d$ case is obvious. In $2d$ two extrema at $\varepsilon=\pm 2$ yield the discontinuities of $D_{2}(\varepsilon)$, while at the saddle point $\varepsilon=0$ the DOS is logarithmically divergent $D_{2}(\varepsilon)\propto-\ln\varepsilon^{2}$ [35]. The both types of singularities fall into the case $\gamma=0$, as understood in the theory of phase transitions. Similarly, calculating the area $\mathcal{S}_{2}(\varepsilon)$ from the quadratic expansion of the spectrum $\varepsilon(\mathbf{k})$ near $\varepsilon=\mp 2$, ($s=0,2$, two terms of the same sign) is sufficient to recover the exact result (79) to leading order: $\mathcal{P}_{2}\propto|\varepsilon\pm 2|$ near the band ends. The quadratic expansion of $\varepsilon(\mathbf{q})$ near the saddle (van Hove) point ($s=1$, two terms of the opposite sign) agrees to leading order with the exact result (80): $\mathcal{P}_{2}-1\propto-\varepsilon\ln\varepsilon^{2}$. So, we infer $\beta=1$ and the dynamical critical index $z=2$. Similar exercise with the expansions of $\varepsilon(\mathbf{q})$ near the critical points yields the $3d$ FST indices (2), in agreement with those obtained from the exact results, see Sec. A.3. Thus, we proved that the whole class universality FST (2) is determined uniquely by $s$ and $d$, where $d=1,2,3$.
To apply the Morse theory to the HK model, we need to make an extension of the above analysis to account for the occurrence of several bands. For our purposes it suffices to deal with two bands defined on two tori, say left/right (L/R). We extend the definition (91) as follows:
$\forall~\mathbf{k}\in M_{i}^{\varepsilon}\subseteq\mathds{T}_{i}^{d}:~\varepsilon_{i}(\mathbf{k})\leq\varepsilon,~~i=L,R~.$ (101)
(Since in the case of the HK model the bands are distinguished by the constant energy shift only, one can visualize the problem as filling of two communicating tori placed at different ground levels). The resulting manifold:
$M^{\varepsilon}\coloneqq M_{L}^{\varepsilon}\cup M_{R}^{\varepsilon}\subseteq\mathds{T}_{L}^{d}\cup\mathds{T}_{R}^{d}~.$ (102)
As follows from the above definition:
$\mathcal{S}_{d}(M^{\varepsilon})=\mathcal{S}_{d}(M_{L}^{\varepsilon})+\mathcal{S}_{d}(M_{R}^{\varepsilon})~,$ (103)
and
$\chi_{E}(M^{\varepsilon})=\chi_{E}(M_{L}^{\varepsilon})+\chi_{E}(M_{R}^{\varepsilon})~.$ (104)
B.2 Generalizations and universality
Now we want to check how robust the results of the previous subsection are with respect to modifications of the spectrum (Morse function). To compare apples to apples, we preserve periodicity of the spectrum $\varepsilon(\mathbf{k})$ defined on $\mathds{T}^{d}$. For instance we can include the next-nearest hopping $t^{\prime}>0$ along main axes. The spectrum
$\varepsilon(\mathbf{k})=-\sum_{i=1}^{d}(\cos k_{i}+t^{\prime}\cos(2k_{i}))$ (105)
at $t^{\prime}<1/4$ yields the same critical points and the critical behavior of the case $t^{\prime}=0$, up to irrelevant change of some numbers. At $t^{\prime}>1/4$ an extra IC critical point (or points, depending on $d$) appears, and additional analysis is warranted. In case of $d=1$ the maximum $\varepsilon(\pi)=1-t^{\prime}$ becomes minimum when $t^{\prime}>1/4$. It is a new Lifshitz point where the Fermi sea becomes disconnected. See Fig. 11. The band gets fully filled at the energy $\varepsilon(x_{\scriptscriptstyle IC})=t^{\prime}+1/8t^{\prime}$ corresponding to two maxima at $x_{\scriptscriptstyle IC}=\pm\arccos(-1/4t^{\prime})$. [^11] There are certain subtleties when the band gets filled at the IC points, and not at the edge or at the center. For details, see analysis of the modulated ladder in [1].
Figure 11: (a): circle $\mathds{T}^{1}=\mathds{S}^{1}$ at $t^{\prime}<1/4$, two critical points with indices (97); (b): Deformed circle at $t^{\prime}>1/4$ with four critical points: min (1), $s=0$, $C_{0}=1$; min (2), $s=0$, $C_{0}=1$; two maxima (3,4), $s=1$, $C_{1}=2$.
The Euler characteristic (compare to (98)) at $t^{\prime}>1/4$ is found as follows:
[…]
The cases $d=2,3$ can be readily analyzed along these lines, revealing additional minima, maxima, and saddle points at $t^{\prime}>1/4$.
It is quite pedagogical to demonstrate in detail what happens to the $2d$ punctured doughnut (see Fig. 8) with modifications of the Morse function (67). Leaving the case (105) as an exercise, we present analysis of another interesting extension of (67) in $2d$ with additional hopping ($t_{1}>0$) along diagonals of the elementary plaquette included. The spectrum is [54]:
$\varepsilon(k_{x},k_{y})=-\cos k_{x}-\cos k_{y}-2t_{1}\cos k_{x}\cos k_{y}~.$ (107)
When $t_{1}<1/2$ the critical behavior is essentially the same as in the case $t^{\prime}=0$. At different energy levels the filled area of the torus $M^{\varepsilon}$ follows the path $a\rightarrow b\rightarrow c$ in Fig. 12, i.e., the same transformations as discussed in the context of Fig. 8. At low filling above the minimum $\varepsilon(0,0)=-2-2t_{1}$, $M^{\varepsilon}$ is a cup, see Fig. 12 a. At the saddle (van Hove) critical point $\varepsilon(0,\pi)=\varepsilon(\pi,0)=2t_{1}$ the cup gets two handles attached, and at $\varepsilon>2t_{1}$, $M^{\varepsilon}$ is a punctured (umbilical) doughnut with a hole on the top (see Fig. 12 c), until $M^{\varepsilon}\to\mathds{T}^{2}$ at the maximum $\varepsilon=\varepsilon(\pi,\pi)=2-2t_{1}$. While going through these critical points, the Euler characteristic changes exactly as in Eq. (95).
When $t_{1}>1/2$ the number of the critical points and their locations within the BZ change as follows:
$\mathrm{Min~1}$ $:$ $~~\varepsilon(0,0)=-2-2t_{1},~~s=0,~C_{0}=1,$
$\mathrm{Min~2}$ $:$ $~~\varepsilon(\pi,\pi)=2-2t_{1},~~s=0,~C_{0}=1,$
$\mathrm{S}$ $:$ $~~\varepsilon(\pm x_{\scriptscriptstyle IC},\pm y_{\scriptscriptstyle IC})=1/2t_{1},~~s=1,~C_{1}=4,$
$\mathrm{Max}$ $:$ $~~\varepsilon(0,\pi)=\varepsilon(\pi,0)=2t_{1},~~s=2,~C_{2}=2,$ (108)
where $x_{\scriptscriptstyle IC}=y_{\scriptscriptstyle IC}=\arccos(-1/2t_{1})$. Evolution of the FS within the BZ is shown in Fig. 12 a,d,e,f. The manifold $M^{\varepsilon}$ undergoes quite different transformations along the path $a\rightarrow d\rightarrow e\rightarrow f$ in Fig. 12, until the BZ $\pi,\pi]^{2}\leftrightarrow\mathds{T}^{2}$ gets fully filled. At the second minimum (Lifshitz point) $M^{\varepsilon}$ becomes disconnected (it gets the second cup), and then at the van Hove critical point (four IC saddle points at $\varepsilon=1/2t_{1}$, see Fig. 12 e) four handles are attached to the two cups. $M^{\varepsilon}$ becomes connected again: it can be viewed as a “distorted” umbilical torus with two holes. Its cross-section with two holes at the energy level $1/2t_{1}<\varepsilon<2t_{1}$ (shown in magenta) and two maxima (black dots) at $\varepsilon=2t_{1}$, is schematically drawn in Fig. 12 f. Two membranes are added when the Morse function reaches the maxima, closing $M^{\varepsilon}\mapsto\mathds{T}^{2}$. The Euler characteristic changes at the critical points as follows:
[…]
From the above analysis of the perturbations (105), (107) of the simplest tight-binding spectrum (67) we can draw now some important general conclusions. Microscopic details of the Hamiltonian (e.g., lattice type, couplings, hopping, interactions) do change the number, the order of appearance, and the multiplicity of the critical points of the Morse function (it is the energy function of the conduction band from the physical point of view). In $d=1$ only maxima or minima are possible, while in higher dimensions saddle points occur in addition. Among the non-degenerate critical points the maxima and minima correspond to the continuous band, Mott, or Lifshitz transitions, while the saddle points correspond to the van Hove transitions. The indices $s$ of the critical points lie in the range:
$\mathrm{For~a~given~}d:~~0\leq s\leq d~.$ (110)
The types (97), (94), and (99) exhaust the possibilities, thus expansions near such critical points (band, Mott, Lifshitz, van Hove) will always yield $\beta$, $\gamma$, and $z$ within the FST universality class (2). Along with the standard attributes of the theory of continuous phase transitions, like the order parameter ($\mathcal{P}$), compressibility ($\chi$), etc, characterized by exponents (2), at each critical point the topology of $M^{\varepsilon}$ undergoes a discrete transformation: $n$-cell (or cells) are added, which is signalled by a change of the Euler characteristic. The sequence of those transformations (and $\chi_{E}(M^{\varepsilon})$) depends on the model’s details (cf., e.g., Figs. 8,11, and 12), but the totally filled band corresponds to $M^{\varepsilon}\mapsto\mathds{T}^{2}$, so $\chi_{E}=0$ for the band and Mott insulators.
The above analysis is not applicable for the degenerate critical points where the Hessian of $\varepsilon(\mathbf{k})$ vanishes. For example, the critical points of functions (105), (107) are degenerate at $t^{\prime}=1/4$ and $t_{1}=1/2$, respectively. Such high-order critical points occur only at some particular (fine-tuned) values of the microscopic parameters, and we assume the situation with non-degenerate critical points of the Morse function as generic. The critical exponents at the high-order points are different from (2) and need to be studied on the case-by-case basis. See, e.g., [49, 50, 51] and more references there.
Appendix C Special points in the M2-phase (d=2)
In case $U<2$ the order of the Lifshitz and the van Hove transitions change with the respect to the case shown in Fig. 5 (c). The van Hove transition occurs inside the M2-phase with two FSs. See Fig. 13. In case of two FSs (M2-metal), we get from the LT (55), (32) the exact result:
$\frac{d\mu}{d\mu_{\circ}}=\frac{2\mathbf{K}(1-\mu_{\circ}^{2}/4)}{\mathbf{K}(1-\mu_{+}^{2}/4)+\mathbf{K}(1-\mu_{-}^{2}/4)}~,$ (111)
which becomes
$\frac{d\mu}{d\mu_{\circ}}=\frac{2\mathbf{K}(1-\mu_{\circ}^{2}/4)}{\mathbf{K}(1-\mu_{\pm}^{2}/4)}$ (112)
in the M-phase with only one FS. From comparison of Eqs. (111) and (112) we infer immediately occurrence of a discontinuity in the slope of $\mu(\mu_{\circ})$ at the Lifshitz points separating the M2- and M-phases.
Figure 13: $d=2$, $U<2$ ($U_{c}=4$). The compressibility $\chi(\mu)$ for $U=1.6$. The critical values of chemical potential: $\pm\mu_{c}=\pm 2.8$, $\pm\mu_{L}=\pm 1.2$. The van Hove critical points correspond $\pm\mu_{vH}=\pm 0.8$ ($\pm\mu_{\circ,vH}=\pm 0.5898$ ) and $\bar{n}(-\mu_{\circ,vH})=0.283547$, $\bar{n}(\mu_{\circ,vH})=0.716438$. The special case $U=2$ when the Lifshitz and van Hove critical points merge is shown in the inset. The Euler characteristics $\chi_{E}$ for each phase are shown in parentheses.
At the van Hove critical points $\mu=\pm U/2$ with arbitrary $U$ the LT (55) yields
$2g(\pm\mu_{\circ})=\pm g_{R}(U)~,$ (113)
and the filling is
$\bar{n}=\frac{1}{2}\pm\frac{1}{8\pi^{2}}g_{R}(U)~.$ (114)
As folows from Eqs. (113) and (114), for $U>2$, when $g_{R}(U)=2\pi^{2}$, the transition occurs at commensurate fillings $\bar{n}=1/4,3/4$, corresponding to the values of noninteracting chemical potential $\mp\mu_{\circ,s}=\mp 0.7198$, cf. Eq. (81):
$g(\mu_{\circ,s})=\pi^{2}~\mapsto~n(\mp\mu_{\circ,s})=1/4,3/4~.$ (115)
This case is analyzed in the main text, Sec. III.3. For $U<2$ the van Hove transition occurs at a generic IC filling
$\bar{n}(\pm\mu_{\circ,vH})=\frac{1}{2}\pm\frac{1}{8\pi^{2}}g(U)~,$ (116)
where the critical values of the noninteracting chemical potential $\mu_{\circ}=\pm\mu_{\circ,vH}$ are found from the following equation:
$g(\mu_{\circ,vH})=\frac{1}{2}g(U)~.$ (117)
This is easy to understand qualitatively from comparison of Figs. 5 (c) and 13: in case of larger separation of the upper and lower bands ($U>2$), the lower band becomes 1/2-filled at $\mu_{vH,1}=-\mu_{vH}$ when the upper band is still empty, while at $\mu_{vH,2}=\mu_{vH}$ the lower band is full and the upper band gets 1/2-filled. The separation of the bands is smaller at $U<2$, and 1/2-filling of the lower band at $\mu_{vH,1}$ occurs simultaneously with some partial filling of the upper band (lower $\leftrightarrow$ upper at $\mu_{vH,2}$).
Using in Eq. (55) the expansions (82) and (80) near the critical points, we get to lowest order the same equation for $\mu(\mu_{\circ},U)$ as Eq. (60), with the only difference that on its l.h.s. $\mu_{\circ,s}\mapsto\mu_{\circ,vH}$ and $g_{1}\coloneqq g^{\prime}(\mu_{\circ,s})\mapsto g^{\prime}(\mu_{\circ,vH})$. Note that at the van Hove points only one of the terms $\mathbf{K}(1-\mu_{\pm}^{2}/4)$ in the denominator of (111) is logarithmically divergent, while the rest is regular. So, the results for $\mu(\mu_{\circ})$ near the van Hove transition in the M-phase, stated around Eqs. (60) and (61), hold for the M2-phase, up to inessential differences in the numbers $g^{\prime}$ ($\propto$ DOS) and $\mu_{\circ,vH}$. Note also that conclusions in the main text about the change of the FS topological properties and of the Euler number at the van Hove points, are not affected wether the latter occur in the M-phase ($U>2$) at special commensurate fillings or in the M2-phase ($U<2$) at rational fillings. In both cases the van Hove transition corresponds to the saddle point at the center of one the two tori, see Appendix B. For completeness we indicated the Euler characteristic for each phase in Fig. 13.
Analysis of the special point $U=2$, where the Lifshitz and van Hove critical points merge (see inset in Fig. 13), shows that the leading order solution (61) for $\mu(\mu_{\circ})$ is applicable in this case with $\mu_{\circ,vH}\to\mu_{\circ,\scriptscriptstyle L}\to\mu_{\circ,\scriptscriptstyle s}$, cf. Eqs. (117), (63), (115), and (81). The occurrence of the van Hove/Lifshitz transitions at either special commensurate, or generic incommensurate fillings does not contradict the plateau theorem, since they take place in the gapless phases.
Data Availability
The additional data that support the findings of this article are available from the author upon reasonable request.